OpenStudy (anonymous):

@Uri @bibby Can you answer this? Choose the slope-intercept equation of the line that passes through the point (2, -3) and is perpendicular to y = two thirdsx + 3. Choose one answer. a. y = -three halvesx b. y = three halvesx - 6 c. y = -two thirdsx - 5 thirds d. y = two thirdsx - 13 thirds

OpenStudy (bibby):

what is the slope of a line perpendicular to \( y = \frac{2}{3}x + 3\)?

OpenStudy (anonymous):

wat do u mean?

OpenStudy (bibby):

what is the slope of the line we want?

OpenStudy (anonymous):

3 is slope.

OpenStudy (uri):

lmao.

OpenStudy (bibby):

y=mx+b where m is the slope

OpenStudy (bibby):

ily uri

OpenStudy (anonymous):

2/3

OpenStudy (uri):

^

OpenStudy (bibby):

perpendicular lines have negative reciprocal slopes of the original lines

OpenStudy (anonymous):

dat means?

OpenStudy (bibby):

reciprocals are "flipped". \(\large \frac{6}{5}\longrightarrow\frac{5}{6}\)

OpenStudy (bibby):

negative reciprocals are negated&& flipped 6/5->-5/6

OpenStudy (anonymous):

i think its c.

OpenStudy (bibby):

nope

OpenStudy (bibby):

guess each letter until you exhaust all possibilities

OpenStudy (anonymous):

no not fair to you or myself.

OpenStudy (bibby):

what is the negative reciprocal of \(\frac{2}{3}\)

OpenStudy (anonymous):

3/2?

OpenStudy (bibby):

negative. negate it

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

b?

OpenStudy (bibby):

why not a

OpenStudy (bibby):

to find a line that passes through (2, -3) with a slope of -3/2 \(y-y_1=\frac{-3}{2}(x-x_1)\) \(y--3=\frac{-3}{2}(x-2)\) \(y+3=\frac{-3}{2}(x-2)\) \(\large y+3=\frac{-3}{2}x-\frac{-3}{\cancel{2}}*\cancel {2}\) \(\large y+3=\frac{-3}{2}x+3\) subtract 3 from both sides \(\large y=\frac{-3}{2}x\) subtract 3 from both sides